Question 3: (10 points). (Yield to maturity) The market price is $750 for a 20-year bond ($1,000 par value) that pays 9 percent annual interest, but makes interest payments on a semiannual basis (4.5 percent semiannually). What is the bonds yield to maturity? (Round to two decimal places.) The bonds yield to maturity is | | %
Answer #1

Answer:

Given

Market price P=$750

year to maturity =20 years

Number of periods N=20*2=40 semiannual

Face Value F=$1000

Coupon rate=9% annual

Semi annual coupon rate=4.5%

So semi annual coupon =4.5%*1000=$45

Let r is the yield to maturity so

P=C*(1-(1+r)^-N)/r + F/(1+r)^N

750=45*(1-(1+r)^-40)/r + 1000/(1+r)^40

So solving for r we get

yield to maturity r=6.205%